40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10758     Entry time: Fri Dec 5 02:44:43 2014
Author: Jenne 
Type: Update 
Category: General 
Subject: IFO alignment shenanigans 

[Jenne, Q, Diego]

OMG, today sucked alignment-wise.  Like, wow. 

I think that the problem with the ASS is with the input pointing part of the system.  I found that if I disable the TTs for the Yarm (iin practice, the outputs are held at zero), I could run the Yarm ASS at full gain of 1, and it would do an awesome job.  The first time I did this, I by-hand optimized the TTs by running the test mass loops to make them follow the input pointing.  After that, I haven't touched the TT pointing at all, and we've just been running the test mass loops for the Yarm ASS.  The Xarm seems to not have this problem (or at least not as drastically), so I left it as it was, touching BS as well as ITMX and ITMY, although the gain still needs to be about 0.3.

I feel pretty good about the IFO alignment now, although it is slightly different than it has been.  The transmitted arm powers are higher than they were before I changed the ASS procedure, and there seems to be a little less power fluctuation with alignment.  Q points out that I don't have concrete evidence that this is a good alignment, but it feels right. 

It was a significant enough change that I had to go down to the Yend to realign the green to the new arm axis.  Xgreen we did with the remote PZTs.  I also realigned both of the beatnotes on the PSL table. 

While I was on the PSL table, I quickly touched up the PMC alignment.

The biggest problem, the one that sucked up more hours and energy than I'd like to admit, is ETMX's jumping.  So frustrating.  Sometimes it is time-coincident with engaging the LSC, sometimes not.  I thought that it might be because there are too many violin filters, but even if I turn off all violin filters to ETMX, it jumped a few times while the cavity was locked.  Sometimes it moves when the cavity is just locked and seems happy, sometimes it moves when nothing is resonating except for the green.  It takes a few minutes to recover the alignment enough to lock, and then it'll jump again a few minutes later.  I haven't gone down to squish the cables today, although I did it yesterday and that didn't seem to do anything. 

We had a few hours of time when it wasn't jumping, so we tried a few times to lock the IFO.  The last several times we have lost lock because the PRC loop rang up.  We measured the loop at low-ish arm powers, but it kept losing lock at higher powers before we could measure.  At least 3 times, the PRC lockloss took out CARM and DARM too.

Anyhow, it has been a long day of not accomplishing anything interesting, but hopefully the IFO will feel better tomorrow.

ELOG V3.1.3-